### v7

# Deep learning in a bilateral brain with hemispheric specialization

Abstract
The brains of all bilaterally symmetric animals on Earth are divided into left and right hemispheres. The anatomy and functionality of the hemispheres have a large degree of overlap, but there are asymmetries and they specialize to possess different attributes. Several computation models mimic hemispheric asymmetries with a focus on reproducing human data on semantic and visual processing tasks. In this study, we aimed to understand how dual hemispheres in a bilateral architecture interact to perform well in a given task. We propose a bilateral artificial neural network that imitates lateralization observed in nature: that the left hemisphere specializes in specificity and the right in generalities. We used different training objectives to achieve the desired specialization and tested it on an image classification task with two different CNN backbones – ResNet and VGG. Our analysis found that the hemispheres represent complementary features that are exploited by a network head which implements a type of weighted attention. The bilateral architecture outperformed a range of baselines of similar representational capacity that don’t exploit differential specialization, with the exception of a conventional ensemble of unilateral networks trained on a dual training objective for specifics and generalities. The results demonstrate the efficacy of bilateralism, contribute to the discussion of bilateralism in biological brains and the principle may serves as an inductive bias for new AI systems.
Impact statement A striking feature of the brain is that it is divided into two hemispheres. It is a conserved feature across all bilaterally symmetric animals, strongly suggesting its importance for intelligence. Many researchers have investigated observed hemispheric asymmetry. We take a different approach, assuming specialization, which enables an investigation into how dual hemispheres complement each other to improve performance. The study constitutes a novel approach to understanding the role of dual hemispheres for the cognitive sciences and the analysis suggests possible mechanisms. In addition, the architecture outperformed a variety of comparable baselines, demonstrating that bilateralism is an important biologically-inspired principle to be considered in the design of new AI systems.
Keywords Hemispheric specialization, Hemispheric asymmetry, Brain-inspired architecture, Bilateral, Computational cognitive neuroscience, Deep learning, Inductive bias
1 Introduction
Many advances in the field of AI have been inspired by the human brain including the perceptron, reinforcement learning and imitation learning. In this study, we took inspiration from the bilateral nature of the brain and applied it to a standard task suited to a bilateral architecture.
What is bilateralism? The brain is divided into left and right hemispheres. It’s a remarkably conserved feature across species, suggesting its importance for intelligence. The anatomy and functionality of the hemispheres have a large degree of overlap, but they are asymmetric and specialize to possess different attributes.
A likely explanation for the emergent specialization is hemispheric asymmetry through small differences in functional and anatomical properties of the cortex. For example, neuron density distribution. The left resembles a pyramid with a greater density of neurons near input layers, whereas the right resembles an inverted pyramid, with greater neuron density near output layers [1, 2, 3, 4, 5, 6]. Network topology also plays a role. The left has more short local pathways constituting a modular network organization, whereas the right has more longer inter-regional pathways [4, 5, 6]. More differences are described in Section 2.1.1, in the Related Work.
Many researchers have used computational models to study the bilateral brain. They mimic anatomical and functional asymmetries, subject the system to standard tasks, and compare the results to human behavioral data. The various studies provide evidence that asymmetric functional and anatomical differences can give rise to observed lateralization of activity (i.e. specialization) for specific tasks. More detail is given in Section 2.1.2, in the Related Work.
The central question of the existing models, is to understand what gives rise to observed specialization. The aim of this project is to investigate a different question – what is the benefit of having differentially specialized hemispheres and how do they complement each other for a given task? Our approach is to induce specialization directly, rather than test if it emerges as a result of architectural asymmetry. We investigated bilateralism through a computational model, with the objective of developing a theory of operation and to understand how bilateral principles could be exploited for AI/ML.
1.1 Our approach
As a starting point, we focused on the left hemisphere’s specialization on specifics, and the right’s specialization on generalities [4, 5, 6]. We hypothesize that an ANN with differentially spec ialized sub-networks would outperform a single comparable network in a classification task featuring both fine and coarse class labels.
We chose a hierarchical image classification task where each sample belongs to both a fine and a coarse class. For example, coarse: sea creature, fine: penguin, seal, shark etc. We modelled the hemispheres with left and right CNNs and induced specialization using supervised training with different training objectives, using either coarse or fine labels. The bilateral network was compared to several comparative baselines that do not have specialization but do have the same number of learnable parameters. The differences were analyzed to explore the effects of bilateral specialization. Experiments were repeated across two types of CNN hemispheres, VGG and ResNet variants, to test generalization over different core CNN architecture.
We predicted that with specialization, the hemispheres would extract distinct but overlapping features, despite identical observations. In addition, we expected that a bilateral architecture would be able to combine these features to improve overall accuracy.
The source code is available at [7].
2 Related work
2.1 Biological computational models
Multiple studies examine hemispheric function in standard behavioral tasks. The studies use bilateral artificial neural networks that mimic functional and architectural hemispheric asymmetries. They replicate aspects of human behavioral data on the tasks, focussing on lateralized hemispheric activity. Some of the studies also consider neurological development, disorders or damage and recovery. With the exception of face perception studies, they all used synthetic abstracted representations for input and output vectors and do not operate on realistic sensory data.
This section is divided into two subsections: a) ‘Hemispheric asymmetries’ which provides additional background on observed architectural and anatomical asymmetries, and b) ‘Models and tasks’ which describes the models that were created, organized by the tasks that were used for testing them.
2.1.1 Hemispheric asymmetries
In addition to the examples given in the introduction, topology and neuron density distribution, other architectural differences come into play. From a connectivity perspective, there are different intra-hemispheric patterns of connectivity between regions within each hemisphere [8, 9]. Furthermore, inter-hemispheric communication by the Corpus collosum results in a competitive process between the hemispheres [2, 1, 10, 11] which is believed to support lateralization of activity. Asymmetry in the visual stream is caused by a split-fovea, where each hemisphere receives the contra-lateral half of the visual field [12, 13, 14, 11, 15, 16, 17].
There are also functional asymmetries. Neurotransmitters such as dopamine and norepinephrine are asymmetrically distributed and neuron firing thresholds differ between hemispheres [10]. Also, the hemispheres respond asymmetrically to frequency: the left responds to higher frequencies than the right [14, 15, 12]. Finally, there may be additional asymmetries that are not yet discovered.
2.1.2 Models and tasks
The behavioral tasks can be divided into two major groups, semantic processing and visual perception. For semantic processing, behavioral data and patterns of hemispheric asymmetry were reproduced by varying intra-hemispheric connectivity [9] in the context of semantic dementia [18], and in the context of unilateral and bilateral lesions [8]. Other studies varied anatomical features such as inter-hemispheric connectivity and the distribution of units between layers, as well as functional features such as learning rate, sensitivity and maximum unit activation and combinations of these factors [2, 1, 10] and in the context of lesions and recovery [3]. Another anatomical feature that was explored is the split-fovea [12] and in the context of dyslexia [13]. In the case of the split-fovea experiments, hemispheric asymmetry emerged as a result of the orthographic to semantic patterns inherent in the language.
Visual perception tasks can be further divided into face perception and the line-bisection task. Face perception includes identification of faces and emotions. In the line-bisection tasks, subjects are asked to mark the center of a horizontal line, as perception of the center is affected by hemispheric function. Face perception was first studied by Dailey et al. Dailey and Cottrell [15] with a model that mimicked the hemispheric frequency response asymmetry. Hsiao et al. Hsiao et al. [14] added a split-fovea and then Wang et al. Wang and Cottrell [11] incorporated other functional parameter differences between the hemispheres as well as a learned gating mechanism for inter-hemispheric communication. For the line bisection task, Shillcock et al. Shillcock and Monaghan [16] and Monaghan and Shillcock Monaghan and Shillcock [17] used a split-fovea, asymmetric frequency response as well as asymmetric distribution of units in layers between left and right.
In the closest work to this project, Mayan et al. Mayan et al. [19] tackled the same problem of image classification with fine and coarse (hierarchical) classes. The authors created a single network with two hemispheres, and trained it with supervised learning and a single loss function. They used hyperparameters with analogies to biological parameters in each hemisphere to encourage specialization. Specialization was achieved in each side but the bilateral network did not have an advantage over baselines.
2.2 Biologically inspired multi-network architectures
Several studies cover related ideas, although they did not explicitly model bilateral hemispheres. Beaulieu et al. [20] created a dual-stream architecture called Neuromodulator Meta-learner, where one network learns to modulate the other, to enhance continual learning. Bakhtiari et al. [21] created a network with parallel pathways to reproduce the functionality of dorsal (‘where’) and ventral (‘what’) pathways in an ANN trained with a single loss function. Li and Deza [22] discussed specialization of a branched neural network when trained on a Gabor filter dataset and found that the training curriculum is inconsequential to specialization of the branched neural network.
2.3 Ensemble models
The dual-hemisphere architecture in this work can be viewed as a type of ensemble model. Ensembling is a popular approach in ML [23]. The most common techniques involve combining the outputs of the same model type, trained with different seeds. However, there are also many techniques that encourage diversity within the ensemble using different training objectives, sampling, architecture, or losses [24, 25, 26]. Our work may be seen as a special case of a diversified ensemble, where the ensemble architecture exploits hierarchical data labels to achieve the diversification.
2.4 Multi-domain learning
Another related area is multi-domain learning. In multi-domain learning (MDL), multiple models are trained to specialize on a particular domain [27]. Two of the main approaches include a) modelling domain-specific and domain-general features [28] and b) explicitly modelling the relationship between domains. MDL is focussed on classes from domains with different distributions. The domain is sometimes explicit, and other times it is determined by the model [29]. In contrast, this project is focussed on explicit hierarchical labels from the same distribution. The objective being to mimic learning different aspects of a single task with left and right hemispheres (2 separate models).
2.5 Hierarchical image datasets
We focused on image classification of fine and coarse classes, exploiting the hierarchical nature of CIFAR-100 [30]. CIFAR-100 and other hierarchical image classification datasets are widely used as benchmarks, including ImageNet [31] and Omniglot [32]. However, the hierarchical nature of available classes is rarely exploited to enhance training or generate a more difficult task. Notable exceptions arise in the Continual Few-Shot Learning literature, such as the Meta-Dataset of Triantafillou et al. [33] which generates class-hierarchy-aware episodes of data and the Tiered-ImageNet of Caccia et al. [34] which uses class hierarchy to generate Out of Distribution (OoD) data.
3 Models
We implemented a bilateral model as well as several baselines. Each model consists of one or more feature extractors, representing hemispheres, as well as one or two classifier heads. The hemispheres were implemented with two different CNN architectures to test if the effects generalize across architecture types: ResNet [35] and VGG [36]. We chose VGG and ResNet because they are common, well understood and amongst the best performing on many vision tasks.
In the case of the ResNet-based models, we took the concept of residual layers and skip connections, and empirically optimized the number of layers on the dataset, resulting in a 9-layer network shown in Figure 2, and hereafter referred to as ResNet-9.
For the VGG based model, we chose the VGG-11A variant [36], hereafter referred to as VGG-11, as it is closest in depth to ResNet-9 and has been shown to work well for CIFAR100 [37]. We used the pre-defined PyTorch implementation, but removed the final fully-connected layer as we later add our own classifier (explained below). In shorthand, the full architecture is: conv64-pool-conv128-pool-conv256-conv256-pool-conv512-conv512-pool-conv512-conv512-pool-fc4096-fc4096. ConvX = convolutional with X filters, pool = max pool, fcX = fully connected with X units. All filters are 3x3 pixels. Although VGG-11 and ResNet-9 are both convolutional networks with a similar depth, there are substantial architectural differences. VGG does not have skip connections, and there are a lot more parameters due to the fully-connected layers.
The heads consist of a single fully-connected layer, without bias or non-linearity and with output dimensions the same size as the number of classes (10 for coarse and 100 for fine).
All networks were trained with a cross-entropy loss function, which includes a softmax operation (not shown explicitly in the diagrams as it is part of the loss function). Details about training and evaluation are given in Section 4.1.2.
The models are depicted in Figure 1 and described in the subsections below.
Figure 1: The range of models. Feature extractors are either ResNet-9 or VGG-11. The classifier heads are single layer fully connected networks, with output dimensions the same size as the number of classes. For brevity, if not specified, the network is unspecialized.
3.1 Bilateral network - specialized
Training and assembling the bilateral network consisted of two phases. In Phase 1, individual hemispheres were trained to be specialized feature extractors with a single classifier head, Figure 1a, referred to as ‘Unilateral - specialized’. The left hemisphere was specialized for specific classes through supervised training using the fine labels. The right hemisphere was specialized for general classes through supervised training using the coarse labels. See Section 4.1.2 for details on training and evaluation.
Then in Phase 2, the hemispheres were brought together into a broader architecture consisting of two hemispheres, which serve as feature extractors, and two newly initialized classifier heads, one each for fine and coarse labels, Figure 1b, referred to as ‘Bilateral - specialized’. The weights of the individual hemispheres were frozen and the heads were trained to use the features for classifying fine and coarse labels. A single loss was calculated by adding the loss from each head.
Figure 2: ResNet-9 architecture. A convolutional neural net with 9 layers and skip connections. The format of the description within each convolutional layer is filter w x h, number of filters, stride and padding. BN = Batch Norm, MP = Max Pool and AP = Average Pool.
3.2 Baseline models
Unilateral network with specialization
To understand the effect of specialization on performance, we compared to the specialized single-hemispheres, Figure 1a, as the first baseline. Specialization was achieved using supervised training with selective labels as described above in Section 3.1, ‘Bilateral network - specialized’.
Unilateral network without specialization
To better understand the single-hemisphere, and the role of using complementary loss functions in the absence of dual hemispheres, we also compared to a single-hemisphere without specialization, Figure 1d. The single hemisphere was trained on both types of labels simultaneously, by using two classifier heads and adding the loss from both.
Bilateral network without specialization
To better understand the role of specialization, we compared the bilateral network to an equivalent network without specialization, Figure 1c. This baseline has the same number of trainable parameters. It can be viewed as a type of ensemble, where the models are combined in the same way as the specialized bilateral network.
We trained the entire network (two hemispheres and two heads) without first explicitly inducing specialization in the individual hemispheres. Like for the unilateral network and bilateral network with specialization, the loss from each head are added together to create a single loss value.
Ensembles
The bilateral network is a type of ensemble. Ensemble learning denotes combining different models or algorithms to obtain improved performance over a single model [23]. To understand the differences between differential specialization and conventional ensembling, we compared to a 2-model ensemble, Figure 1e. We chose 2 models, so that the ensemble has approximately the same number of trainable parameters as the bilateral network. We used two unilateral models without specialization, trained from different seeds. The output of the ensemble (in training and inference) was the average output from the heads: for the fine classification, it was the average of the fine classifier heads, and for the coarse classification, it was the average of the coarse classifier heads. In order to test multiple seeds, we trained five models with different seeds, and then constructed five 2-model ensembles, each with a unique pair of models selected from the pool of five.
4 Experiments
4.1 Experimental setup
4.1.1 Dataset
We used the CIFAR-100 dataset [30], as it includes hierarchical labels that denote fine (specific) and coarse (general) class labels. The dataset is split into 50,000 training and 10,000 test samples. Images were resized to 32x32 and we used two widely adopted data augmentation techniques, a random crop (using the RandomCrop PyTorch function) and a random horizontal flip (using the RandomHorizontalFlip PyTorch function) to increase the diversity of the training set.
4.1.2 Training and evaluation
The models were trained with supervised learning using a cross-entropy loss function. Depending on the model, the fine or coarse labels were used. In the case of dual heads (one with each label type, see Section 3), the loss from each head was added to create a single loss value. Weights were initialized with PyTorch defaults, from a random uniform distribution $U(-\sqrt{k},\sqrt{k})$: for linear layers $k=1/in\_features$, and for the convolutional layers $k=\frac{1}{C_{in}*\Pi_{i=0}^{1}kernel\_size[i]}$. We used weight decay of 1.0e-5, dropout of 0.6 on the classifier heads [38] to regularize the models, Adam [39] for optimization, with a learning rate of 1.0e-4 and a mini-batch size of 256.
Each network was trained for 180 epochs from 5 random seeds. At the end of each epoch, the models were validated on the test split. The number of epochs was sufficient to ensure that the validation accuracy had plateaued, and the final model was the one with the highest validation accuracy across epochs.
4.1.3 Framework and computational resources
Experiments were conducted with the PyTorch-Lightning [40] research framework, a wrapper around the PyTorch library [41]. All models were trained and evaluated with a virtual machine with a single NVIDIA A10 GPU, 24GB RAM (through Lambda Labs https://lambdalabs.com/).
4.2 Visualizing network operation
For visualizations, we focussed on one hemisphere architecture, ResNet-9. We used two types of visualizations to better understand the operation of the bilateral network.
In order to be informative, we selected key scenarios that are distinct from each other and highlight the contribution of different parts of the network. The scenarios are:
  • Scenario 1: The bilateral-specialized network is correct, left and right hemispheres are incorrect.
  • Scenario 2: The bilateral-specialized network and the right hemisphere are correct, the left hemisphere is incorrect.
  • Scenario 3: The bilateral-specialized network and the left hemisphere are correct, the right hemisphere is incorrect.
  • Scenario 4: The bilateral-specialized network is incorrect, the left and right hemispheres are correct.
The definition of ‘correct’ for the bilateral-specialized network, is that it was successful for both fine and coarse class labels.
Gradient camera (Grad-Cam) visualization
To understand how the specialized bilateral network exploits features from both hemispheres, we visualized gradient flow with respect to target convolutional layers [42, 43, 44] using the Grad-Cam library. The gradients reveal which areas of the image contribute most to the prediction; the gradient heatmap highlights the region of focus. We applied the visualization to each of the specialized hemispheres individually, and compared it to the bilateral network.
For a single hemisphere, the gradients were averaged over the 2nd convolutional layer of each residual block (2 convolutional layers with skip connections, see Figure 2) during prediction. In the case of the bilateral network, the gradients were averaged over both hemispheres during simultaneous prediction of both coarse and fine labels (the network had two heads).
Feature analysis using cosine similarity
We investigated how left and right features are exploited by the bilateral network by analyzing the relationship between representations in different parts of the network, with a focus on how the representations are transformed by the network heads. We did this by measuring the similarity of features for images of the same label. They are expected to have similar features, so measuring the feature similarity at different parts of the network should be revealing.
We first grouped the image samples into random pairs with the same class label. We then plotted a bivariate distribution of cosine similarity between the pairs, one for the input to the heads, denoted ‘concatenated’ and the other for the average of the network head outputs, denoted ‘bilateral’. Similarity is plotted along the following dimensions: left hemisphere, right hemisphere and a third dimension, either concatenated or bilateral. Additionally, univariate marginal distributions were plotted for each hemisphere.
5 Results
5.1 Accuracy
Quantitative results are shown in tabular form in Tables 1 and 2 and with plots in Figures 3 and 4.
The relative performance of models was very consistent between ResNet and VGG architectures. In summary, the bilateral architecture with specialization provided a performance boost for both fine and coarse classes over an individual specialized hemisphere (unilateral-specialized). Moreover, it outperformed the baselines, except for the 2-model ensemble. The 2-model ensemble consists of unilateral models, which were more effective than unilateral that were specialized in most cases. The exception was the VGG architecture with coarse labels, where the unspecialized unilateral model was not better than the specialized version, and the resulting 2-model ensemble did not outperform the specialized bilateral network.
Model # Params (M)
Accuracy (%)
- fine labels
Accuracy (%)
- coarse labels
Unilateral 6.63 67.81 $\pm$ 0.46 78.25 $\pm$ 0.51
Unilateral (specialized for fine) 6.58 67.81 $\pm$ 0.31
Unilateral (specialized for coarse) 6.62 75.87 $\pm$ 0.70
Bilateral 13.26 68.80 $\pm$ 0.59 78.82 $\pm$ 0.65
Bilateral - with specialization 13.26 71.35 $\pm$ 0.36 80.71 $\pm$ 0.48
Ensemble (2 unilateral models) 13.26 72.15 $\pm$ 0.11 81.80 $\pm$ 0.24
Table 1: Accuracy of ResNet-based models. Accuracy is given as mean $\pm$ 1 standard deviation of 5 seeds.
Model # Params (M)
Accuracy (%)
- fine labels
Accuracy (%)
- coarse labels
Unilateral 129.26 53.13 $\pm$ 0.41 66.86 $\pm$ 0.41
Unilateral (specialized for fine) 129.18 51.87 $\pm$ 0.58
Unilateral (specialized for coarse) 128.85 66.86 $\pm$ 0.41
Bilateral 258.52 53.05 $\pm$ 0.49 66.42 $\pm$ 0.27
Bilateral with specialization 258.52 56.09 $\pm$ 0.33 71.57 $\pm$ 0.42
Ensemble (2 unilateral models) 258.52 57.81 $\pm$ 0.38 70.68 $\pm$ 0.30
Table 2: Accuracy of VGG-based models. Accuracy is given as mean $\pm$ 1 standard deviation of 5 seeds.
Figure 3: Accuracy of ResNet-based architectures. For brevity, if not specified, the network is unspecialized. The height of the bars is the mean accuracy, and the error bars are 1 standard deviation of 5 seeds.
Figure 4: Accuracy of VGG-11-based models. For brevity, if not specified, the network is unspecialized. The height of the bars is the mean accuracy, and the error bars are 1 standard deviation of 5 seeds.
5.2 Visualizing network operation
Visualizations of Grad-Cam and similarity distributions are shown for the selected scenarios described in Section 4.2.
The Grad-Cam visualizations are shown in Figures 5 to 8. In general, the features are more local in the left hemisphere and more global in the right hemisphere. The heads blend different aspects of features from the left and right hemispheres for the task. The effect is that in many cases, features trained for fine classes are helpful for coarse classes and vice versa, and a correct classification can be achieved even if both hemispheres are individually incorrect.
(a) Bosc Image
(b) Left hemisphere
(c) Right hemisphere
(d) Bilateral-specialized
Figure 5: Grad-Cam for Scenario 1: The bilateral-specialized network is correct, left and right hemispheres are incorrect. The bilateral network appears to adjust the area of focus even when both the hemispheres’ features are situated external to the bosc.
(a) Tank Image
(b) Left hemisphere
(c) Right hemisphere
(d) Bilateral-specialized
Figure 6: Grad-Cam for Scenario 2: The bilateral-specialized network and the right hemisphere are correct, the left hemisphere is incorrect. Possibly due to the fact that there are multiple labels with wheels, the left failed to differentiate the tank by its wheels alone, however the bilateral appears to overcome this, using the right hemisphere’s features.
(a) Cycle Image
(b) Left hemisphere
(c) Right hemisphere
(d) Bilateral-specialized
Figure 7: Grad-Cam for Scenario 3: The bilateral-specialized network and the left hemisphere are correct, the right hemisphere is incorrect. The left hemisphere identified more local features of the bicycle compared to the right. The network head appear to exploit the left representations to make an accurate prediction.
(a) Sofa Image
(b) Left hemisphere
(c) Right hemisphere
(d) Bilateral-specialized
Figure 8: Grad-Cam for Scenario 4: The bilateral-specialized network is incorrect, the left and right hemispheres are correct. Both hemispheres identified local and global features. The network heads appear to over-compensate and focus on the region outside of the sofa, which is ultimately unsuccessful.
The cosine similarity results are shown in Figure 9 and Figure10. In the concatenated distribution, there is a strong correlation between similarity in left and concatenated, and right and concatenated. In contrast, there is no obvious correlation in the bilateral distribution. The network heads learned a non-linear transformation of the feature space.
(a) Concatenated
(b) Bilateral-specialized network
Figure 9: Cosine similarity distribution for Scenario 1: The bilateral-specialized network is correct, left and right hemispheres are incorrect. Many of the pairs (same label) have dissimilar (values closer to 0) features in left and right, seen in the density of points close to the origin. The concatenated features are also dissimilar, but the bilateral network elevates the similarity of many of the points. Each point is a pair of images with the same label. The x-axis is similarity in the left hemisphere, y-axis is similarity in the right hemisphere. The color is similarity in the combined representation. Additionally, the univariate marginal distributions are shown for left and right with histograms above and to the side of the bivariate distribution.
(a) Concatenated
(b) Bilateral-specialized network
Figure 10: Cosine similarity distribution for Scenario 2: The bilateral-specialized network and the right hemisphere are correct, the left hemisphere is incorrect. Some regions with points that have low similarity in the left (at approximately x, y = 0.2, 0.2), have elevated similarity in the bilateral output, compared to concatenated. Also, for points with high similarity in the left (approximately 0.35, 0.1), the similarity is lower in the bilateral output. In other words, the bilateral network was able to exploit features in the right hemisphere over the left. Each point is a pair of images with the same label. The x-axis is similarity in the left hemisphere, y-axis is similarity in the right hemisphere. The color is similarity in the combined representation. Additionally, the univariate marginal distributions are shown for left and right with histograms above and to the side of the bivariate distribution.
6 Discussion
6.1 Key findings
We conducted experiments to implement and study bilateralism in Artificial Neural Networks. The effects of bilateralism were studied by comparing against baselines that captured distinct characteristics of the proposed bilateral network. We used a classification task with hierarchical classes that captured an observed characteristic of biological hemispheres, that the left is more specialized for fine classes and the right for coarse classes.
We found that there is an advantage of two hemispheres over one, demonstrated by the fact that Bilateral-specialized outperformed individual hemispheres alone. Furthermore, we found that it is not due to the greater number of learnable parameters, and having two hemispheres is not sufficient, but that specialization is important, shown by the benefit of bilateral-specialized over bilateral.
Surprisingly, an unspecialized hemisphere (Unilateral) performed as well or better than specialized hemispheres (Unilateral-specialized) in most cases. Training with two objective functions (fine and coarse) led to better performance on both labels – whether that is in a bilateral or unilateral architecture.
Ultimately, two hemispheres arranged into a conventional ensemble achieved the best results. A key difference with Bilateral-specialized, is that the component models were more effective to begin with. Another difference is the way in which the model outputs are combined – averaging of the classification vs concatenation of the feature vector. Future work could explore this in more detail.
Both the specialized bilateral network and the ensemble were able to boost performance of individual hemispheres. However, they use different mechanisms.
The novel finding of the paper is that a bilateral architecture can exploit specialized hemispheres, and the remainder of the paper examines this principles.
6.2 Computational cost
Compared to a conventional single network, the bilateral architecture has approximately double the number of parameters and computational cost. The performance boost between unilateral and specialized bilateral models may appear modest for the additional cost. However, incremental gains are difficult to achieve at high levels of performance, making the  2.5-5% improvement (across fine, coarse and hemisphere architectures) consequential. This is the first study to demonstrate the bilateral principle, and like other new architectures, greater improvements can be expected in the future.
6.3 How does the bilateral architecture work?
The Grad-Cam images reveal that the left hemisphere extracts more localized features than the right. Different learning objectives enable them to capture different aspects of the environment. Collectively the set of features is greater than one network with one objective.
Interestingly, even though the left is explicitly trained on fine class labels, the features that it extracts are helpful for coarse classes. The inverse is true of the right.
With reference to the cosine similarity visualization, before the network heads, the similarity of pairs of images of the same class is correlated with left and right. The similarity is no longer correlated after the network heads, which must implement some sort of non-linear transformation. In many cases, when left or right produce ineffective features, shown by low similarity between these images of the same class, the heads are able to produce features that have increased similarity. The network heads learn to combine the features from left and right, to different degrees, to produce better predictions. In some cases, the bilateral network makes a correct classification, even if both hemispheres are individually wrong.
In summary, specialization creates a higher diversity of features. The network heads implement a type of weighted attention to left and right hemispheres selectively in a task dependent manner, improving overall class prediction.
6.4 Relation to biology
A limitation of our model, in terms of biological accuracy, was that inter-hemispheric connectivity was achieved at the outputs of the hemispheres. In contrast, biological hemispheres are interconnected throughout their hierarchies [45]. Nevertheless, the heads may serve a similar purpose, albeit in a cruder way. Inter-hemispheric connectivity comprises a complex combination of inhibitory and excitatory projections. The hemisphere that is better able to represent the input is likely to have stronger activation and thus inhibit the other hemisphere [1, 2]. Like the heads, this too is a type of selectivity. Building a model with interconnected hemispheres throughout the hierarchy is a topic for future research.
The central finding that in our model, left and right hemispheres extracted local and non-local features, used together in fine tasks, provides one plausible suggestion for the prevalence of bilateral brains in nature. A possible approach to investigate this idea is with transcranial magnetic stimulation (TMS) to selectively impair one hemisphere at a time on controlled tasks [46].
6.5 Why it’s important
The key findings are a small step towards understanding and taking advantage of the characteristics of biological brains, and show the potential of bilateralism to improve ML/AI. More broadly, it’s a biological principle that warrants further investigation.
The fact that bilateralism has a material effect on the network as a whole, suggests that bilateralism should be considered for other cognitive neuroscience models, where it is usually ignored. One potentially fruitful area is extending the standard Hippocampal/Neocortical model, CLS [47, 48], by introducing asymmetries into the bilateral Hippocampal architecture. Indeed, evidence exists of asymmetric contributions of left and right Hippocampi to memory [49, 50].
Bilateralism could also inspire new priors or inductive biases to improve ML models.
6.6 Future work
Given that conventional ensembling was effective, and in addition differential specialization provided benefits, it would be interesting to examine the advantage of each. For example, does one of these methods take you further on different or more difficult datasets? Furthermore, are there ways to combine aspects of both methods, such ensembling bilateral models – would this be more effective than simply increasing the size of the ensemble for the same total number of hemispheres. Finally, can we leverage the specialized features in more effective ways, such as using a combiner network as opposed to simple concatenation of features.
A promising direction to improve the model as well as our understanding of the neuroscience, is to further investigate and model the neurobiology. There are several avenues. For example, replicating recurrent connectivity, more complex biologically inspired interactions between the hemispheres, mimicking known substrate differences between the hemispheres such as topological differences, resource allocation (see Figure LABEL:fig:bilateral_figure) and experiments on inducing specialization without supervision.
This study focused on one aspect of bilateralism with an image classification task. There are also interesting lateral effects for motor control. For example, in most people, the right hand is better at trajectories, and the left hand is better at position/posture (dynamic dominance theory [51]). Developing an architecture and experiments to explore motor control is an interesting area for future work.
One theory is that the observed differences between the hemispheres emerge from a more fundamental specialization for novelty (right) and routine (left) [4, 5, 52]. These ideas could be explored with an agent that can act in a dynamic environment. The right hemisphere could be a generalist that can perform unfamiliar tasks as a beginner, while the left becomes an expert over time. The agent would be able to adopt new tasks, without being inept at them. Currently, the field of Continual RL does not focus on avoiding poor performance, but rather maximizing the best performance. In real-life scenarios however, an agent must avoid death and serious injury to themselves and those around them (also relevant for physical robots and virtual artificial agents).
As mentioned earlier in this section, bilateralism could be incorporated into existing cognitive computational models, and the benefits of bilateralism could be studied within more powerful ML architectures.
7 Conclusion
We built a bilateral architecture with left and right neural networks, inspired by our bilateral brains. The hemispheres were trained to specialize on fine and coarse classes. They extracted specialized features and through a type of ‘weighted attention’ by a simple fully connected layer, outperformed various (but not all) baselines on classification of both fine and coarse classes. The specialized representations had benefits above the explicit objective of their individual hemispheres. The results demonstrate that small procedural changes to training can achieve specialization, and that specialization can be complementary and beneficial for certain tasks. The operation of the artificial network provides ideas for the study of neurobiological hemispheric specialization. Simultaneously, this work shows that neuroscientific principles can provide inductive biases for novel ML architectures. Currently, in the field of AI/ML where scaling of existing architectures is achieving great success, it’s interesting to look at new principles on smaller architecture that could then be scaled.
Acknowledgments
In an earlier Masters project, co-supervised by Levin Kuhlmann, Amir Mayan explored similar ideas, which formed a valuable background to planning this project. Thanks to Punarjay Chakravarty for helpful discussions.
Author contributions
GK conceived and supervised the project, CR designed and implemented the experiments, GK and CR authored the manuscript, DR advised and assisted with the manuscript and EG introduced the concept and provided advice. All authors read and approved the final manuscript.

### v8

@@
 The brains of all bilaterally symmetric animals on Earth are divided into left and right hemispheres.
-The anatomy and functionality of the hemispheres have a large degree of overlap, but there are asymmetries and they specialize to possess different attributes.
-Several computation models mimic hemispheric asymmetries with a focus on reproducing human data on semantic and visual processing tasks.
-In this study, we aimed to understand how dual hemispheres in a bilateral architecture interact to perform well in a given task.
-We propose a bilateral artificial neural network that imitates lateralization observed in nature: that the left hemisphere specializes in specificity and the right in generalities.
-We used different training objectives to achieve the desired specialization and tested it on an image classification task with two different CNN backbones – ResNet and VGG.
+The anatomy and functionality of the hemispheres have a large degree of overlap, but there are asymmetries and they specialise to possess different attributes.
+Other authors have used computational models to mimic hemispheric asymmetries with a focus on reproducing human data on semantic and visual processing tasks.
+We took a different approach and aimed to understand how dual hemispheres in a bilateral architecture interact to perform well in a given task.
+We propose a bilateral artificial neural network that imitates lateralisation observed in nature: that the left hemisphere specialises in specificity and the right in generality.
+We used different training objectives to achieve the desired specialisation and tested it on an image classification task with two different CNN backbones – ResNet and VGG.
 Our analysis found that the hemispheres represent complementary features that are exploited by a network head which implements a type of weighted attention.
-The bilateral architecture outperformed a range of baselines of similar representational capacity that don’t exploit differential specialization, with the exception of a conventional ensemble of unilateral networks trained on a dual training objective for specifics and generalities.
+The bilateral architecture outperformed a range of baselines of similar representational capacity that don’t exploit differential specialisation, with the exception of a conventional ensemble of unilateral networks trained on a dual training objective for specifics and generalities.
 The results demonstrate the efficacy of bilateralism, contribute to the discussion of bilateralism in biological brains and the principle may serves as an inductive bias for new AI systems.
-Impact statement A striking feature of the brain is that it is divided into two hemispheres.
-It is a conserved feature across all bilaterally symmetric animals, strongly suggesting its importance for intelligence.
-Many researchers have investigated observed hemispheric asymmetry.
-We take a different approach, assuming specialization, which enables an investigation into how dual hemispheres complement each other to improve performance.
-The study constitutes a novel approach to understanding the role of dual hemispheres for the cognitive sciences and the analysis suggests possible mechanisms.
-In addition, the architecture outperformed a variety of comparable baselines, demonstrating that bilateralism is an important biologically-inspired principle to be considered in the design of new AI systems.
-Keywords Hemispheric specialization, Hemispheric asymmetry, Brain-inspired architecture, Bilateral, Computational cognitive neuroscience, Deep learning, Inductive bias
+Keywords Bilateral  $\cdot$ Hemispheric specialisation  $\cdot$ Hemispheric asymmetry  $\cdot$ Brain-inspired architecture  $\cdot$ Cognitive architectures  $\cdot$ Deep learning  $\cdot$ Computational cognitive neuroscience  $\cdot$ Inductive bias
 1 Introduction
-Many advances in the field of AI have been inspired by the human brain including the perceptron, reinforcement learning and imitation learning.
-In this study, we took inspiration from the bilateral nature of the brain and applied it to a standard task suited to a bilateral architecture.
-What is bilateralism?
-The brain is divided into left and right hemispheres.
-It’s a remarkably conserved feature across species, suggesting its importance for intelligence.
-The anatomy and functionality of the hemispheres have a large degree of overlap, but they are asymmetric and specialize to possess different attributes.
-A likely explanation for the emergent specialization is hemispheric asymmetry through small differences in functional and anatomical properties of the cortex.
-For example, neuron density distribution.
-The left resembles a pyramid with a greater density of neurons near input layers, whereas the right resembles an inverted pyramid, with greater neuron density near output layers [1, 2, 3, 4, 5, 6].
-Network topology also plays a role.
-The left has more short local pathways constituting a modular network organization, whereas the right has more longer inter-regional pathways [4, 5, 6].
-More differences are described in Section 2.1.1, in the Related Work.
-Many researchers have used computational models to study the bilateral brain.
-They mimic anatomical and functional asymmetries, subject the system to standard tasks, and compare the results to human behavioral data.
-The various studies provide evidence that asymmetric functional and anatomical differences can give rise to observed lateralization of activity (i.e.
-specialization) for specific tasks.
-More detail is given in Section 2.1.2, in the Related Work.
-The central question of the existing models, is to understand what gives rise to observed specialization.
-The aim of this project is to investigate a different question – what is the benefit of having differentially specialized hemispheres and how do they complement each other for a given task?
-Our approach is to induce specialization directly, rather than test if it emerges as a result of architectural asymmetry.
-We investigated bilateralism through a computational model, with the objective of developing a theory of operation and to understand how bilateral principles could be exploited for AI/ML.
+Left and right hemispheres are a remarkably conserved feature of brains across species, suggesting their importance for intelligence.
+The anatomy and functionality of the hemispheres have a large degree of overlap, but they are asymmetric and specialise to possess different attributes.
+For example, the left hemisphere is specialised for specific (fine) classes and familiarity/routine whereas the right hemisphere is specialised for more general (coarse) classes and novelty [4, 5].
+Researchers have studied the bilateral brain with computational models that mimic anatomical and functional asymmetries.
+The models are typically subjected to standard tasks and compared to human behavioural data.
+The various studies provide evidence that asymmetric functional and anatomical differences can give rise to observed lateralisation of activity (i.e.
+specialisation) for specific tasks e.g.
+[1, 18, 13, 8, 11].
+The central question of existing models is to understand the cause of the observed specialisation.
+The aim of this project is to investigate a different question – what is the benefit of specialised hemispheres and how do they complement each other for a given task?
+Our approach is to induce specialisation directly, rather than test if it emerges as a result of architectural asymmetry.
+Our objective is to develop a theory of operation and to understand how bilateral principles could be exploited for AI/ML.
 1.1 Our approach
-As a starting point, we focused on the left hemisphere’s specialization on specifics, and the right’s specialization on generalities [4, 5, 6].
-We hypothesize that an ANN with differentially spec ialized sub-networks would outperform a single comparable network in a classification task featuring both fine and coarse class labels.
-We chose a hierarchical image classification task where each sample belongs to both a fine and a coarse class.
-For example, coarse: sea creature, fine: penguin, seal, shark etc.
-We modelled the hemispheres with left and right CNNs and induced specialization using supervised training with different training objectives, using either coarse or fine labels.
-The bilateral network was compared to several comparative baselines that do not have specialization but do have the same number of learnable parameters.
-The differences were analyzed to explore the effects of bilateral specialization.
-Experiments were repeated across two types of CNN hemispheres, VGG and ResNet variants, to test generalization over different core CNN architecture.
-We predicted that with specialization, the hemispheres would extract distinct but overlapping features, despite identical observations.
+We focused on the left hemisphere’s specialisation on specifics, and the right’s specialisation on generalities [4, 5].
+We hypothesised that an ANN with differentially specialised sub-networks would outperform a single comparable network in a classification task featuring both fine and coarse class labels.
+We predicted that with specialisation, the hemispheres would extract distinct but overlapping features, despite identical observations.
 In addition, we expected that a bilateral architecture would be able to combine these features to improve overall accuracy.
+We chose a hierarchical image classification task where each sample belongs to both a fine and a coarse class e.g.
+coarse: sea creature, fine: penguin, seal, shark.
+We modelled the hemispheres with left and right CNNs and induced specialisation using supervised training with different objectives, using either coarse or fine labels.
+The bilateral network was compared to several comparative baselines that do not have specialisation but do have the same number of learnable parameters.
+The differences were analysed to explore the effects of bilateral specialisation.
+Experiments were repeated with two types of CNN hemispheres, VGG and ResNet variants, to test generalisation over different backbones.
 The source code is available at [7].
 2 Related work
-2.1 Biological computational models
-Multiple studies examine hemispheric function in standard behavioral tasks.
-The studies use bilateral artificial neural networks that mimic functional and architectural hemispheric asymmetries.
-They replicate aspects of human behavioral data on the tasks, focussing on lateralized hemispheric activity.
-Some of the studies also consider neurological development, disorders or damage and recovery.
-With the exception of face perception studies, they all used synthetic abstracted representations for input and output vectors and do not operate on realistic sensory data.
-This section is divided into two subsections: a) ‘Hemispheric asymmetries’ which provides additional background on observed architectural and anatomical asymmetries, and b) ‘Models and tasks’ which describes the models that were created, organized by the tasks that were used for testing them.
-2.1.1 Hemispheric asymmetries
-In addition to the examples given in the introduction, topology and neuron density distribution, other architectural differences come into play.
-From a connectivity perspective, there are different intra-hemispheric patterns of connectivity between regions within each hemisphere [8, 9].
-Furthermore, inter-hemispheric communication by the Corpus collosum results in a competitive process between the hemispheres [2, 1, 10, 11] which is believed to support lateralization of activity.
-Asymmetry in the visual stream is caused by a split-fovea, where each hemisphere receives the contra-lateral half of the visual field [12, 13, 14, 11, 15, 16, 17].
-There are also functional asymmetries.
-Neurotransmitters such as dopamine and norepinephrine are asymmetrically distributed and neuron firing thresholds differ between hemispheres [10].
-Also, the hemispheres respond asymmetrically to frequency: the left responds to higher frequencies than the right [14, 15, 12].
-Finally, there may be additional asymmetries that are not yet discovered.
-2.1.2 Models and tasks
-The behavioral tasks can be divided into two major groups, semantic processing and visual perception.
-For semantic processing, behavioral data and patterns of hemispheric asymmetry were reproduced by varying intra-hemispheric connectivity [9] in the context of semantic dementia [18], and in the context of unilateral and bilateral lesions [8].
-Other studies varied anatomical features such as inter-hemispheric connectivity and the distribution of units between layers, as well as functional features such as learning rate, sensitivity and maximum unit activation and combinations of these factors [2, 1, 10] and in the context of lesions and recovery [3].
-Another anatomical feature that was explored is the split-fovea [12] and in the context of dyslexia [13].
-In the case of the split-fovea experiments, hemispheric asymmetry emerged as a result of the orthographic to semantic patterns inherent in the language.
-Visual perception tasks can be further divided into face perception and the line-bisection task.
-Face perception includes identification of faces and emotions.
-In the line-bisection tasks, subjects are asked to mark the center of a horizontal line, as perception of the center is affected by hemispheric function.
-Face perception was first studied by Dailey et al.
-Dailey and Cottrell [15] with a model that mimicked the hemispheric frequency response asymmetry.
-Hsiao et al.
-Hsiao et al.
-[14] added a split-fovea and then Wang et al.
-Wang and Cottrell [11] incorporated other functional parameter differences between the hemispheres as well as a learned gating mechanism for inter-hemispheric communication.
-For the line bisection task, Shillcock et al.
-Shillcock and Monaghan [16] and Monaghan and Shillcock Monaghan and Shillcock [17] used a split-fovea, asymmetric frequency response as well as asymmetric distribution of units in layers between left and right.
-In the closest work to this project, Mayan et al.
-Mayan et al.
-[19] tackled the same problem of image classification with fine and coarse (hierarchical) classes.
-The authors created a single network with two hemispheres, and trained it with supervised learning and a single loss function.
-They used hyperparameters with analogies to biological parameters in each hemisphere to encourage specialization.
-Specialization was achieved in each side but the bilateral network did not have an advantage over baselines.
-2.2 Biologically inspired multi-network architectures
-Several studies cover related ideas, although they did not explicitly model bilateral hemispheres.
+Biologically inspired multi-network architectures
 Beaulieu et al.
@@
 [21] created a network with parallel pathways to reproduce the functionality of dorsal (‘where’) and ventral (‘what’) pathways in an ANN trained with a single loss function.
-Li and Deza [22] discussed specialization of a branched neural network when trained on a Gabor filter dataset and found that the training curriculum is inconsequential to specialization of the branched neural network.
-2.3 Ensemble models
-The dual-hemisphere architecture in this work can be viewed as a type of ensemble model.
-Ensembling is a popular approach in ML [23].
-The most common techniques involve combining the outputs of the same model type, trained with different seeds.
-However, there are also many techniques that encourage diversity within the ensemble using different training objectives, sampling, architecture, or losses [24, 25, 26].
-Our work may be seen as a special case of a diversified ensemble, where the ensemble architecture exploits hierarchical data labels to achieve the diversification.
-2.4 Multi-domain learning
-Another related area is multi-domain learning.
-In multi-domain learning (MDL), multiple models are trained to specialize on a particular domain [27].
+Li and Deza [22] discussed specialisation of a branched neural network.
+In the most related study, Mayan et al.
+[19] tackled the same image classification task with fine and coarse (hierarchical) classes.
+The authors created a single network with two hemispheres, and trained it with supervised learning and a single loss function.
+They used hyperparameters with analogies to biological parameters in each hemisphere to encourage specialisation.
+Specialisation was achieved in each side but the bilateral network did not have an advantage over baselines.
+Ensemble models
+Our bilateral approach can be viewed as a type of ensemble, which is a popular approach in ML [23].
+Typically, the outputs of multiple models, trained on different seeds, are combined; and the models are of the same type.
+Other techniques encourage diversity within the ensemble using different training objectives, sampling, architecture, or losses [24, 25, 26].
+Our work can be viewed as a special case of a diversified ensemble, where the ensemble architecture exploits hierarchical data labels to achieve diversification.
+Multi-domain learning
+In multi-domain learning (MDL), multiple models are trained to specialise in a particular domain [27].
 Two of the main approaches include a) modelling domain-specific and domain-general features [28] and b) explicitly modelling the relationship between domains.
-MDL is focussed on classes from domains with different distributions.
-The domain is sometimes explicit, and other times it is determined by the model [29].
-In contrast, this project is focussed on explicit hierarchical labels from the same distribution.
+MDL is focused on classes from domains with different distributions.
+The domain is sometimes explicit and other times it is determined by the model [29].
+In contrast, this project is focused on explicit hierarchical labels from the same distribution.
 The objective being to mimic learning different aspects of a single task with left and right hemispheres (2 separate models).
-2.5 Hierarchical image datasets
-We focused on image classification of fine and coarse classes, exploiting the hierarchical nature of CIFAR-100 [30].
-CIFAR-100 and other hierarchical image classification datasets are widely used as benchmarks, including ImageNet [31] and Omniglot [32].
-However, the hierarchical nature of available classes is rarely exploited to enhance training or generate a more difficult task.
-Notable exceptions arise in the Continual Few-Shot Learning literature, such as the Meta-Dataset of Triantafillou et al.
-[33] which generates class-hierarchy-aware episodes of data and the Tiered-ImageNet of Caccia et al.
-[34] which uses class hierarchy to generate Out of Distribution (OoD) data.
 3 Models
-We implemented a bilateral model as well as several baselines.
-Each model consists of one or more feature extractors, representing hemispheres, as well as one or two classifier heads.
-The hemispheres were implemented with two different CNN architectures to test if the effects generalize across architecture types: ResNet [35] and VGG [36].
+We compared a bilateral model to several baselines.
+Each model consisted of one or more feature extractor, representing hemispheres, as well as one or two classifier heads, one for each label type (fine or coarse).
+The hemispheres were implemented with two different CNN architectures to test if the effects generalise across architectures: ResNet [35] and VGG [36].
 We chose VGG and ResNet because they are common, well understood and amongst the best performing on many vision tasks.
-In the case of the ResNet-based models, we took the concept of residual layers and skip connections, and empirically optimized the number of layers on the dataset, resulting in a 9-layer network shown in Figure 2, and hereafter referred to as ResNet-9.
-For the VGG based model, we chose the VGG-11A variant [36], hereafter referred to as VGG-11, as it is closest in depth to ResNet-9 and has been shown to work well for CIFAR100 [37].
-We used the pre-defined PyTorch implementation, but removed the final fully-connected layer as we later add our own classifier (explained below).
+In the case of ResNet-based models, we empirically optimised the number of layers on the dataset, resulting in a 9-layer network, hereafter referred to as ResNet-9.
+In shorthand, the full architecture is: conv64-conv128-pool-res128,128-pool-conv256-conv512-pool-res512,512-av_pool.
+ConvX = convolutional with X filters, res = residual block with two convolutional blocks of X,Y filters, pool = max pool, av_pool = average pool.
+All filters are 3x3 and we used ReLU activation function.
+For the VGG model, we chose the VGG-11A variant [36], hereafter referred to as VGG-11, as it is closest in depth to ResNet-9 and works well for CIFAR-100 [37].
+We used the PyTorch implementation, but removed the final fully-connected layer as we added our own classifier (explained below).
 In shorthand, the full architecture is: conv64-pool-conv128-pool-conv256-conv256-pool-conv512-conv512-pool-conv512-conv512-pool-fc4096-fc4096.
-ConvX = convolutional with X filters, pool = max pool, fcX = fully connected with X units.
-All filters are 3x3 pixels.
+Definitions as above, and fcX = fully connected with X units.
+All filters are 3x3.
 Although VGG-11 and ResNet-9 are both convolutional networks with a similar depth, there are substantial architectural differences.
@@
 The heads consist of a single fully-connected layer, without bias or non-linearity and with output dimensions the same size as the number of classes (10 for coarse and 100 for fine).
-All networks were trained with a cross-entropy loss function, which includes a softmax operation (not shown explicitly in the diagrams as it is part of the loss function).
-Details about training and evaluation are given in Section 4.1.2.
-The models are depicted in Figure 1 and described in the subsections below.
+The models are depicted in Figure 1 and described below. Details on training and evaluation in Section 4.1.2.
 Figure 1: The range of models.
@@
 The classifier heads are single layer fully connected networks, with output dimensions the same size as the number of classes.
-For brevity, if not specified, the network is unspecialized.
-3.1 Bilateral network - specialized
-Training and assembling the bilateral network consisted of two phases.
-In Phase 1, individual hemispheres were trained to be specialized feature extractors with a single classifier head, Figure 1a, referred to as ‘Unilateral - specialized’.
-The left hemisphere was specialized for specific classes through supervised training using the fine labels.
-The right hemisphere was specialized for general classes through supervised training using the coarse labels.
-See Section 4.1.2 for details on training and evaluation.
-Then in Phase 2, the hemispheres were brought together into a broader architecture consisting of two hemispheres, which serve as feature extractors, and two newly initialized classifier heads, one each for fine and coarse labels, Figure 1b, referred to as ‘Bilateral - specialized’.
+For brevity, if not specified, the network is unspecialised.
+3.1 Bilateral network - specialised
+Assembling and training the network consisted of two phases.
+In Phase 1, individual hemispheres were trained to be specialised feature extractors using supervised learning with a single classifier head, resulting in ‘Unilateral - specialised’, Figure 1a.
+The left was specialised for specific classes using fine labels, and the right for general classes with coarse labels.
+In Phase 2, the hemispheres (serving as feature extractors) were brought together into a broader architecture, with two initialised classifier heads, one for fine and coarse labels, resulting in ‘Bilateral - specialised’, Figure 1b, .
 The weights of the individual hemispheres were frozen and the heads were trained to use the features for classifying fine and coarse labels.
 A single loss was calculated by adding the loss from each head.
-Figure 2: ResNet-9 architecture.
-A convolutional neural net with 9 layers and skip connections.
-The format of the description within each convolutional layer is filter w x h, number of filters, stride and padding.
-BN = Batch Norm, MP = Max Pool and AP = Average Pool.
 3.2 Baseline models
-Unilateral network with specialization
-To understand the effect of specialization on performance, we compared to the specialized single-hemispheres, Figure 1a, as the first baseline.
-Specialization was achieved using supervised training with selective labels as described above in Section 3.1, ‘Bilateral network - specialized’.
-Unilateral network without specialization
-To better understand the single-hemisphere, and the role of using complementary loss functions in the absence of dual hemispheres, we also compared to a single-hemisphere without specialization, Figure 1d.
-The single hemisphere was trained on both types of labels simultaneously, by using two classifier heads and adding the loss from both.
-Bilateral network without specialization
-To better understand the role of specialization, we compared the bilateral network to an equivalent network without specialization, Figure 1c.
-This baseline has the same number of trainable parameters.
-It can be viewed as a type of ensemble, where the models are combined in the same way as the specialized bilateral network.
-We trained the entire network (two hemispheres and two heads) without first explicitly inducing specialization in the individual hemispheres.
-Like for the unilateral network and bilateral network with specialization, the loss from each head are added together to create a single loss value.
+Unilateral networks with specialisation
+We compared the bilateral network to the individual specialised hemispheres alone, on the tasks for which they were specialised, referred to as Unilateral-specialised, Figure 1a.
+Specialisation was achieved using supervised training with selective labels as described in Section 3.1.
+Unilateral network without specialisation
+To investigate specialisation, we compared Unilateral-specialised to a single hemisphere without specialisation, referred to as Unilateral-unspecialised, Figure 1d.
+The single hemisphere was trained on both types of labels simultaneously, by using two classifier heads and adding the loss.
+This baseline also helps to understand the effect of complementary loss functions; on a single hemisphere compared to the bilateral network.
+Bilateral network without specialisation
+To better understand the role of specialisation in the bilateral network, we compared Bilateral-specialised to an equivalent network, with the same number of trainable parameters, but without specialisation, referred to as Bilateral-unspecialised, Figure 1c.
+It can be viewed as a type of ensemble, where the models are combined in the same way as the specialised bilateral network.
+We trained the entire network (two hemispheres and two heads) without first explicitly inducing specialisation in the individual hemispheres.
+Like for the unilateral network and bilateral network with specialisation, the loss from each head are added together to create a single loss.
 Ensembles
-The bilateral network is a type of ensemble.
-Ensemble learning denotes combining different models or algorithms to obtain improved performance over a single model [23].
-To understand the differences between differential specialization and conventional ensembling, we compared to a 2-model ensemble, Figure 1e.
+To understand the differences between differential specialisation and conventional ensembling, we compared to a 2-model ensemble, Figure 1e.
 We chose 2 models, so that the ensemble has approximately the same number of trainable parameters as the bilateral network.
-We used two unilateral models without specialization, trained from different seeds.
-The output of the ensemble (in training and inference) was the average output from the heads: for the fine classification, it was the average of the fine classifier heads, and for the coarse classification, it was the average of the coarse classifier heads.
-In order to test multiple seeds, we trained five models with different seeds, and then constructed five 2-model ensembles, each with a unique pair of models selected from the pool of five.
+We used two unilateral models without specialisation, trained from different seeds.
+The output of the ensemble was the average from the heads: for fine classification, the average of the fine heads, and for coarse, the average of the coarse heads.
+We trained five models with different seeds, and then constructed five 2-model ensembles, each with a unique pair of models selected from the pool of five.
 4 Experiments
@@
 4.1.1 Dataset
-We used the CIFAR-100 dataset [30], as it includes hierarchical labels that denote fine (specific) and coarse (general) class labels.
+We used the CIFAR-100 dataset [30], as it includes hierarchical labels that denote fine (specific) and coarse (general) classes.
 The dataset is split into 50,000 training and 10,000 test samples.
-Images were resized to 32x32 and we used two widely adopted data augmentation techniques, a random crop (using the RandomCrop PyTorch function) and a random horizontal flip (using the RandomHorizontalFlip PyTorch function) to increase the diversity of the training set.
+Images were resized to 32x32 and we used two widely adopted data augmentation techniques, a random crop and a random horizontal flip (using the inbuilt PyTorch functions) to increase the diversity of the training set.
 4.1.2 Training and evaluation
 The models were trained with supervised learning using a cross-entropy loss function.
-Depending on the model, the fine or coarse labels were used.
 In the case of dual heads (one with each label type, see Section 3), the loss from each head was added to create a single loss value.
-Weights were initialized with PyTorch defaults, from a random uniform distribution $U(-\sqrt{k},\sqrt{k})$: for linear layers $k=1/in\_features$, and for the convolutional layers $k=\frac{1}{C_{in}*\Pi_{i=0}^{1}kernel\_size[i]}$.
-We used weight decay of 1.0e-5, dropout of 0.6 on the classifier heads [38] to regularize the models, Adam [39] for optimization, with a learning rate of 1.0e-4 and a mini-batch size of 256.
+Weights were initialised with PyTorch defaults, from a random uniform distribution $U(-\sqrt{k},\sqrt{k})$: for linear layers $k=1/in\_features$, and for the convolutional layers $k=\frac{1}{C_{in}*\Pi_{i=0}^{1}kernel\_size[i]}$.
+We used weight decay of 1.0e-5, dropout of 0.6 on the classifier heads to regularise the models, Adam [39] for optimisation, with a learning rate of 1.0e-4 and a mini-batch size of 256.
 Each network was trained for 180 epochs from 5 random seeds.
@@
 All models were trained and evaluated with a virtual machine with a single NVIDIA A10 GPU, 24GB RAM (through Lambda Labs https://lambdalabs.com/).
-4.2 Visualizing network operation
-For visualizations, we focussed on one hemisphere architecture, ResNet-9. We used two types of visualizations to better understand the operation of the bilateral network.
-In order to be informative, we selected key scenarios that are distinct from each other and highlight the contribution of different parts of the network. The scenarios are:
-  • Scenario 1: The bilateral-specialized network is correct, left and right hemispheres are incorrect.
-  • Scenario 2: The bilateral-specialized network and the right hemisphere are correct, the left hemisphere is incorrect.
-  • Scenario 3: The bilateral-specialized network and the left hemisphere are correct, the right hemisphere is incorrect.
-  • Scenario 4: The bilateral-specialized network is incorrect, the left and right hemispheres are correct.
-The definition of ‘correct’ for the bilateral-specialized network, is that it was successful for both fine and coarse class labels.
-Gradient camera (Grad-Cam) visualization
-To understand how the specialized bilateral network exploits features from both hemispheres, we visualized gradient flow with respect to target convolutional layers [42, 43, 44] using the Grad-Cam library.
+4.2 Visualising network operation
+We focused on one backbone, ResNet-9, and used two types of visualisations to better understand the operation of the bilateral network.
+To be informative, we selected key scenarios that are distinct from each other and highlight the contribution of different parts of the network, described below.
+The definition of ‘correct’ for the bilateral-specialised network, is that it was successful for both fine and coarse class labels.
+  • Scenario 1: Bilateral-specialised network is correct, left and right hemispheres are incorrect
+  • Scenario 2: Bilateral-specialised network and right hemisphere are correct, left hemisphere is incorrect
+  • Scenario 3: Bilateral-specialised network and left hemisphere are correct, right hemisphere is incorrect
+  • Scenario 4: Bilateral-specialised network is incorrect, left and right hemispheres are correct
+Gradient camera (Grad-Cam) visualisation
+To understand how the specialised bilateral network exploits features from both hemispheres, we visualised gradient flow with respect to target convolutional layers [43, 42] using the Grad-Cam library [44].
 The gradients reveal which areas of the image contribute most to the prediction; the gradient heatmap highlights the region of focus.
-We applied the visualization to each of the specialized hemispheres individually, and compared it to the bilateral network.
-For a single hemisphere, the gradients were averaged over the 2nd convolutional layer of each residual block (2 convolutional layers with skip connections, see Figure 2) during prediction.
+We applied the visualisation to each of the specialised hemispheres individually, and compared it to the bilateral network.
+For a single hemisphere, the gradients were averaged over the 2nd convolutional layer of each residual block (2 convolutional layers with skip connections) during prediction.
 In the case of the bilateral network, the gradients were averaged over both hemispheres during simultaneous prediction of both coarse and fine labels (the network had two heads).
 Feature analysis using cosine similarity
-We investigated how left and right features are exploited by the bilateral network by analyzing the relationship between representations in different parts of the network, with a focus on how the representations are transformed by the network heads.
+We investigated how left and right features are exploited by the bilateral network by analysing the relationship between representations in different parts of the network, with a focus on how the representations are transformed by the network heads.
 We did this by measuring the similarity of features for images of the same label.
@@
 5.1 Accuracy
-Quantitative results are shown in tabular form in Tables 1 and 2 and with plots in Figures 3 and 4.
-The relative performance of models was very consistent between ResNet and VGG architectures.
-In summary, the bilateral architecture with specialization provided a performance boost for both fine and coarse classes over an individual specialized hemisphere (unilateral-specialized).
-Moreover, it outperformed the baselines, except for the 2-model ensemble.
-The 2-model ensemble consists of unilateral models, which were more effective than unilateral that were specialized in most cases.
-The exception was the VGG architecture with coarse labels, where the unspecialized unilateral model was not better than the specialized version, and the resulting 2-model ensemble did not outperform the specialized bilateral network.
+The relative performance of models was very consistent between ResNet and VGG backbones, Tables 1 and 2 and Figures 2 and 3.
+Accuracy given as mean $\pm$ 1 standard deviation (shown as error bars in the plots) of 5 seeds.
+In summary, Bilateral-specialised outperformed all the baselines, except for the 2-model ensemble.
+The 2-model ensemble consists of unilateral models, which were more effective than specialised unilateral models in most cases.
+The exception was the VGG backbone with coarse labels, where the unspecialised unilateral model was not better than the specialised version, and the resulting 2-model ensemble did not outperform Bilateral-specialised.
 Model # Params (M)
@@
 Unilateral 6.63 67.81 $\pm$ 0.46 78.25 $\pm$ 0.51
-Unilateral (specialized for fine) 6.58 67.81 $\pm$ 0.31
-Unilateral (specialized for coarse) 6.62 75.87 $\pm$ 0.70
+Unilateral-specialised 6.58 67.81 $\pm$ 0.31
+Unilateral-specialised 6.62 75.87 $\pm$ 0.70
 Bilateral 13.26 68.80 $\pm$ 0.59 78.82 $\pm$ 0.65
-Bilateral - with specialization 13.26 71.35 $\pm$ 0.36 80.71 $\pm$ 0.48
+Bilateral-specialised 13.26 71.35 $\pm$ 0.36 80.71 $\pm$ 0.48
 Ensemble (2 unilateral models) 13.26 72.15 $\pm$ 0.11 81.80 $\pm$ 0.24
-Table 1: Accuracy of ResNet-based models. Accuracy is given as mean $\pm$ 1 standard deviation of 5 seeds.
+Table 1: Accuracy of ResNet-based models.
 Model # Params (M)
@@
 Unilateral 129.26 53.13 $\pm$ 0.41 66.86 $\pm$ 0.41
-Unilateral (specialized for fine) 129.18 51.87 $\pm$ 0.58
-Unilateral (specialized for coarse) 128.85 66.86 $\pm$ 0.41
+Unilateral-specialised 129.18 51.87 $\pm$ 0.58
+Unilateral-specialised 128.85 66.86 $\pm$ 0.41
 Bilateral 258.52 53.05 $\pm$ 0.49 66.42 $\pm$ 0.27
-Bilateral with specialization 258.52 56.09 $\pm$ 0.33 71.57 $\pm$ 0.42
+Bilateral-specialised 258.52 56.09 $\pm$ 0.33 71.57 $\pm$ 0.42
 Ensemble (2 unilateral models) 258.52 57.81 $\pm$ 0.38 70.68 $\pm$ 0.30
-Table 2: Accuracy of VGG-based models. Accuracy is given as mean $\pm$ 1 standard deviation of 5 seeds.
-Figure 3: Accuracy of ResNet-based architectures.
-For brevity, if not specified, the network is unspecialized.
-The height of the bars is the mean accuracy, and the error bars are 1 standard deviation of 5 seeds.
-Figure 4: Accuracy of VGG-11-based models.
-For brevity, if not specified, the network is unspecialized.
-The height of the bars is the mean accuracy, and the error bars are 1 standard deviation of 5 seeds.
-5.2 Visualizing network operation
-Visualizations of Grad-Cam and similarity distributions are shown for the selected scenarios described in Section 4.2.
-The Grad-Cam visualizations are shown in Figures 5 to 8.
-In general, the features are more local in the left hemisphere and more global in the right hemisphere.
-The heads blend different aspects of features from the left and right hemispheres for the task.
+Table 2: Accuracy of VGG-based models
+Figure 2: Accuracy of ResNet-based models. For brevity, if not specified, the network is unspecialised.
+Figure 3: Accuracy of VGG-11-based models. For brevity, if not specified, the network is unspecialised.
+5.2 Visualising network operation
+The Grad-Cam visualisations are shown in Figures 4 to 7 for the scenarios described in Section 4.2.
+In general, the features appear to be more local in the left hemisphere and more global in the right.
+The heads blend different aspects of features from the left and right, for the task.
 The effect is that in many cases, features trained for fine classes are helpful for coarse classes and vice versa, and a correct classification can be achieved even if both hemispheres are individually incorrect.
 (a) Bosc Image
-(b) Left hemisphere
-(c) Right hemisphere
-(d) Bilateral-specialized
-Figure 5: Grad-Cam for Scenario 1: The bilateral-specialized network is correct, left and right hemispheres are incorrect.
+(b) Left
+(c) Right
+(d) Bilateral-sp
+Figure 4: Grad-Cam for Scenario 1: The bilateral-specialised (abbreviated to bilateral-sp in the figure) network is correct, left and right are incorrect.
 The bilateral network appears to adjust the area of focus even when both the hemispheres’ features are situated external to the bosc.
 (a) Tank Image
-(b) Left hemisphere
-(c) Right hemisphere
-(d) Bilateral-specialized
-Figure 6: Grad-Cam for Scenario 2: The bilateral-specialized network and the right hemisphere are correct, the left hemisphere is incorrect.
-Possibly due to the fact that there are multiple labels with wheels, the left failed to differentiate the tank by its wheels alone, however the bilateral appears to overcome this, using the right hemisphere’s features.
+(b) Left
+(c) Right
+(d) Bilateral-sp
+Figure 5: Grad-Cam for Scenario 2: The bilateral-specialised network and the right are correct, the left is incorrect.
+Possibly due to the fact that there are multiple labels with wheels, the left failed to differentiate the tank by its wheels alone, however the bilateral model appears to overcome this, using the right’s features.
 (a) Cycle Image
-(b) Left hemisphere
-(c) Right hemisphere
-(d) Bilateral-specialized
-Figure 7: Grad-Cam for Scenario 3: The bilateral-specialized network and the left hemisphere are correct, the right hemisphere is incorrect.
+(b) Left
+(c) Right
+(d) Bilateral-sp
+Figure 6: Grad-Cam for Scenario 3: The bilateral-specialised network and the left are correct, the right is incorrect.
 The left hemisphere identified more local features of the bicycle compared to the right.
@@
 (a) Sofa Image
-(b) Left hemisphere
-(c) Right hemisphere
-(d) Bilateral-specialized
-Figure 8: Grad-Cam for Scenario 4: The bilateral-specialized network is incorrect, the left and right hemispheres are correct.
+(b) Left
+(c) Right
+(d) Bilateral-sp
+Figure 7: Grad-Cam for Scenario 4: The bilateral-specialised network is incorrect, the left and right hemispheres are correct.
 Both hemispheres identified local and global features.
@@
 In contrast, there is no obvious correlation in the bilateral distribution.
-The network heads learned a non-linear transformation of the feature space.
+The network heads appear to have learned a non-linear transformation of the feature space.
 (a) Concatenated
-(b) Bilateral-specialized network
-Figure 9: Cosine similarity distribution for Scenario 1: The bilateral-specialized network is correct, left and right hemispheres are incorrect.
+(b) Bilateral-specialised network
+Figure 8: Cosine similarity distribution for Scenario 1: The bilateral-specialised network is correct, left and right are incorrect.
 Many of the pairs (same label) have dissimilar (values closer to 0) features in left and right, seen in the density of points close to the origin.
@@
 (a) Concatenated
-(b) Bilateral-specialized network
-Figure 10: Cosine similarity distribution for Scenario 2: The bilateral-specialized network and the right hemisphere are correct, the left hemisphere is incorrect.
+(b) Bilateral-specialised network
+Figure 9: Cosine similarity distribution for Scenario 2: The bilateral-specialised network and the right are correct, the left is incorrect.
 Some regions with points that have low similarity in the left (at approximately x, y = 0.2, 0.2), have elevated similarity in the bilateral output, compared to concatenated.
@@
 6 Discussion
-6.1 Key findings
-We conducted experiments to implement and study bilateralism in Artificial Neural Networks.
-The effects of bilateralism were studied by comparing against baselines that captured distinct characteristics of the proposed bilateral network.
-We used a classification task with hierarchical classes that captured an observed characteristic of biological hemispheres, that the left is more specialized for fine classes and the right for coarse classes.
-We found that there is an advantage of two hemispheres over one, demonstrated by the fact that Bilateral-specialized outperformed individual hemispheres alone.
-Furthermore, we found that it is not due to the greater number of learnable parameters, and having two hemispheres is not sufficient, but that specialization is important, shown by the benefit of bilateral-specialized over bilateral.
-Surprisingly, an unspecialized hemisphere (Unilateral) performed as well or better than specialized hemispheres (Unilateral-specialized) in most cases.
+We found that there is an advantage of two hemispheres over one, demonstrated by the fact that the specialised bilateral architecture (Bilateral-specialised) outperformed individual hemispheres (Unilateral-specialised and Unilateral).
+Furthermore, we found that it is not due to the greater number of learnable parameters, and having two hemispheres is not sufficient, but that specialisation is important, shown by the benefit of Bilateral-specialised over Bilateral.
+Surprisingly, unspecialised hemispheres (Unilateral) performed as well or better than specialised hemispheres (Unilateral-specialised) in most cases.
 Training with two objective functions (fine and coarse) led to better performance on both labels – whether that is in a bilateral or unilateral architecture.
 Ultimately, two hemispheres arranged into a conventional ensemble achieved the best results.
-A key difference with Bilateral-specialized, is that the component models were more effective to begin with.
+A key difference with Bilateral-specialised, is that the component models were more effective to begin with.
 Another difference is the way in which the model outputs are combined – averaging of the classification vs concatenation of the feature vector.
 Future work could explore this in more detail.
-Both the specialized bilateral network and the ensemble were able to boost performance of individual hemispheres. However, they use different mechanisms.
-The novel finding of the paper is that a bilateral architecture can exploit specialized hemispheres, and the remainder of the paper examines this principles.
-6.2 Computational cost
+Both the specialised bilateral network and the ensemble were able to boost performance of individual hemispheres. However, they use different mechanisms.
+The novel finding of the paper is that a bilateral architecture can exploit specialised hemispheres, and the remainder of the paper examines this principle.
+Computational cost
 Compared to a conventional single network, the bilateral architecture has approximately double the number of parameters and computational cost.
-The performance boost between unilateral and specialized bilateral models may appear modest for the additional cost.
+The performance boost between unilateral and specialised bilateral models may appear modest for the additional cost.
 However, incremental gains are difficult to achieve at high levels of performance, making the  2.5-5% improvement (across fine, coarse and hemisphere architectures) consequential.
 This is the first study to demonstrate the bilateral principle, and like other new architectures, greater improvements can be expected in the future.
-6.3 How does the bilateral architecture work?
-The Grad-Cam images reveal that the left hemisphere extracts more localized features than the right.
+How does the bilateral architecture work?
+The Grad-Cam images reveal that the left hemisphere extracts more localised features than the right.
 Different learning objectives enable them to capture different aspects of the environment.
 Collectively the set of features is greater than one network with one objective.
-Interestingly, even though the left is explicitly trained on fine class labels, the features that it extracts are helpful for coarse classes. The inverse is true of the right.
-With reference to the cosine similarity visualization, before the network heads, the similarity of pairs of images of the same class is correlated with left and right.
-The similarity is no longer correlated after the network heads, which must implement some sort of non-linear transformation.
-In many cases, when left or right produce ineffective features, shown by low similarity between these images of the same class, the heads are able to produce features that have increased similarity.
+Interestingly, even though the left is explicitly trained on fine labels, the features that it extracts are helpful for coarse classes.
+The inverse is true of the right.
+With reference to the cosine similarity visualisation, before the network heads, the similarity of pairs of images of the same class is correlated with left and right and the similarity is no longer correlated after the network heads; which suggests that the heads implement a non-linear transformation.
+In many cases, when left or right produce ineffective features, shown by low similarity between embeddings of these images of the same class, the heads are able to produce features that have increased similarity.
 The network heads learn to combine the features from left and right, to different degrees, to produce better predictions.
 In some cases, the bilateral network makes a correct classification, even if both hemispheres are individually wrong.
-In summary, specialization creates a higher diversity of features.
+In summary, specialisation creates a higher diversity of features.
 The network heads implement a type of weighted attention to left and right hemispheres selectively in a task dependent manner, improving overall class prediction.
-6.4 Relation to biology
+Relation to biology
 A limitation of our model, in terms of biological accuracy, was that inter-hemispheric connectivity was achieved at the outputs of the hemispheres.
@@
 Building a model with interconnected hemispheres throughout the hierarchy is a topic for future research.
-The central finding that in our model, left and right hemispheres extracted local and non-local features, used together in fine tasks, provides one plausible suggestion for the prevalence of bilateral brains in nature.
+The central finding, that left and right hemispheres extracted local and non-local features, to improve performance on a given task, provides one plausible suggestion (and testable hypothesis) for the prevalence of bilateral brains in nature.
 A possible approach to investigate this idea is with transcranial magnetic stimulation (TMS) to selectively impair one hemisphere at a time on controlled tasks [46].
-6.5 Why it’s important
-The key findings are a small step towards understanding and taking advantage of the characteristics of biological brains, and show the potential of bilateralism to improve ML/AI.
-More broadly, it’s a biological principle that warrants further investigation.
-The fact that bilateralism has a material effect on the network as a whole, suggests that bilateralism should be considered for other cognitive neuroscience models, where it is usually ignored.
-One potentially fruitful area is extending the standard Hippocampal/Neocortical model, CLS [47, 48], by introducing asymmetries into the bilateral Hippocampal architecture.
-Indeed, evidence exists of asymmetric contributions of left and right Hippocampi to memory [49, 50].
-Bilateralism could also inspire new priors or inductive biases to improve ML models.
-6.6 Future work
-Given that conventional ensembling was effective, and in addition differential specialization provided benefits, it would be interesting to examine the advantage of each.
-For example, does one of these methods take you further on different or more difficult datasets?
-Furthermore, are there ways to combine aspects of both methods, such ensembling bilateral models – would this be more effective than simply increasing the size of the ensemble for the same total number of hemispheres.
-Finally, can we leverage the specialized features in more effective ways, such as using a combiner network as opposed to simple concatenation of features.
-A promising direction to improve the model as well as our understanding of the neuroscience, is to further investigate and model the neurobiology.
-There are several avenues.
-For example, replicating recurrent connectivity, more complex biologically inspired interactions between the hemispheres, mimicking known substrate differences between the hemispheres such as topological differences, resource allocation (see Figure LABEL:fig:bilateral_figure) and experiments on inducing specialization without supervision.
-This study focused on one aspect of bilateralism with an image classification task.
-There are also interesting lateral effects for motor control.
-For example, in most people, the right hand is better at trajectories, and the left hand is better at position/posture (dynamic dominance theory [51]).
-Developing an architecture and experiments to explore motor control is an interesting area for future work.
-One theory is that the observed differences between the hemispheres emerge from a more fundamental specialization for novelty (right) and routine (left) [4, 5, 52].
-These ideas could be explored with an agent that can act in a dynamic environment.
-The right hemisphere could be a generalist that can perform unfamiliar tasks as a beginner, while the left becomes an expert over time.
-The agent would be able to adopt new tasks, without being inept at them.
-Currently, the field of Continual RL does not focus on avoiding poor performance, but rather maximizing the best performance.
-In real-life scenarios however, an agent must avoid death and serious injury to themselves and those around them (also relevant for physical robots and virtual artificial agents).
-As mentioned earlier in this section, bilateralism could be incorporated into existing cognitive computational models, and the benefits of bilateralism could be studied within more powerful ML architectures.
 7 Conclusion
-We built a bilateral architecture with left and right neural networks, inspired by our bilateral brains.
-The hemispheres were trained to specialize on fine and coarse classes.
-They extracted specialized features and through a type of ‘weighted attention’ by a simple fully connected layer, outperformed various (but not all) baselines on classification of both fine and coarse classes.
-The specialized representations had benefits above the explicit objective of their individual hemispheres.
-The results demonstrate that small procedural changes to training can achieve specialization, and that specialization can be complementary and beneficial for certain tasks.
-The operation of the artificial network provides ideas for the study of neurobiological hemispheric specialization.
+Inspired by our bilateral brains, we built a bilateral architecture with left and right neural networks using two CNN backbones and compared them to baselines that captured distinct characteristics of the proposed bilateral network.
+We used a classification task with hierarchical classes that captured an observed characteristic of biological hemispheres, that the left is more specialised for fine classes and the right for coarse classes.
+The hemispheres were trained to specialise on fine and coarse classes accordingly.
+They extracted specialised features and through a type of ‘weighted attention’ by a simple fully connected layer, outperformed various (but not all) baselines on classification of both fine and coarse classes.
+The specialised representations had benefits above the explicit objective of their individual hemispheres.
+The results demonstrate that small procedural changes to training can achieve specialisation, and that specialisation can be complementary and beneficial for certain tasks.
+The operation of the artificial network provides ideas for the study of neurobiological hemispheric specialisation.
 Simultaneously, this work shows that neuroscientific principles can provide inductive biases for novel ML architectures.
 Currently, in the field of AI/ML where scaling of existing architectures is achieving great success, it’s interesting to look at new principles on smaller architecture that could then be scaled.
-Acknowledgments
-In an earlier Masters project, co-supervised by Levin Kuhlmann, Amir Mayan explored similar ideas, which formed a valuable background to planning this project.
+Acknowledgements
 Thanks to Punarjay Chakravarty for helpful discussions.
-Author contributions
-GK conceived and supervised the project, CR designed and implemented the experiments, GK and CR authored the manuscript, DR advised and assisted with the manuscript and EG introduced the concept and provided advice.
-All authors read and approved the final manuscript.

### v9

@@
-Deep learning in a bilateral brain with hemispheric specialization
+https://cerenaut.ai
 Abstract
 The brains of all bilaterally symmetric animals on Earth are divided into left and right hemispheres.
-The anatomy and functionality of the hemispheres have a large degree of overlap, but there are asymmetries and they specialise to possess different attributes.
+The anatomy and functionality of the hemispheres have a large degree of overlap, but there are asymmetries, and they specialise in possesses different attributes.
 Other authors have used computational models to mimic hemispheric asymmetries with a focus on reproducing human data on semantic and visual processing tasks.
 We took a different approach and aimed to understand how dual hemispheres in a bilateral architecture interact to perform well in a given task.
-We propose a bilateral artificial neural network that imitates lateralisation observed in nature: that the left hemisphere specialises in specificity and the right in generality.
-We used different training objectives to achieve the desired specialisation and tested it on an image classification task with two different CNN backbones – ResNet and VGG.
-Our analysis found that the hemispheres represent complementary features that are exploited by a network head which implements a type of weighted attention.
-The bilateral architecture outperformed a range of baselines of similar representational capacity that don’t exploit differential specialisation, with the exception of a conventional ensemble of unilateral networks trained on a dual training objective for specifics and generalities.
-The results demonstrate the efficacy of bilateralism, contribute to the discussion of bilateralism in biological brains and the principle may serves as an inductive bias for new AI systems.
-Keywords Bilateral  $\cdot$ Hemispheric specialisation  $\cdot$ Hemispheric asymmetry  $\cdot$ Brain-inspired architecture  $\cdot$ Cognitive architectures  $\cdot$ Deep learning  $\cdot$ Computational cognitive neuroscience  $\cdot$ Inductive bias
+We propose a bilateral artificial neural network that imitates lateralisation observed in nature: that the left hemisphere specialises in local features and the right in global features.
+We used different training objectives to achieve the desired specialisation and tested it on an image classification task with two different CNN backbones: ResNet and VGG.
+Our analysis found that the hemispheres represent complementary features that are exploited by a network head that implements a type of weighted attention.
+The bilateral architecture outperformed a range of baselines of similar representational capacity that do not exploit differential specialisation, with the exception of a conventional ensemble of unilateral networks trained on dual training objectives for local and global features.
+The results demonstrate the efficacy of bilateralism, contribute to the discussion of bilateralism in biological brains, and the principle may serve as an inductive bias for new AI systems.
+Keywords:
+Hemispheric specialisation Hemispheric asymmetry Brain-inspired architecture Cognitive architectures Inductive bias
 1 Introduction
-Left and right hemispheres are a remarkably conserved feature of brains across species, suggesting their importance for intelligence.
-The anatomy and functionality of the hemispheres have a large degree of overlap, but they are asymmetric and specialise to possess different attributes.
-For example, the left hemisphere is specialised for specific (fine) classes and familiarity/routine whereas the right hemisphere is specialised for more general (coarse) classes and novelty [4, 5].
-Researchers have studied the bilateral brain with computational models that mimic anatomical and functional asymmetries.
-The models are typically subjected to standard tasks and compared to human behavioural data.
-The various studies provide evidence that asymmetric functional and anatomical differences can give rise to observed lateralisation of activity (i.e.
-specialisation) for specific tasks e.g.
-[1, 18, 13, 8, 11].
-The central question of existing models is to understand the cause of the observed specialisation.
-The aim of this project is to investigate a different question – what is the benefit of specialised hemispheres and how do they complement each other for a given task?
-Our approach is to induce specialisation directly, rather than test if it emerges as a result of architectural asymmetry.
-Our objective is to develop a theory of operation and to understand how bilateral principles could be exploited for AI/ML.
-1.1 Our approach
-We focused on the left hemisphere’s specialisation on specifics, and the right’s specialisation on generalities [4, 5].
-We hypothesised that an ANN with differentially specialised sub-networks would outperform a single comparable network in a classification task featuring both fine and coarse class labels.
-We predicted that with specialisation, the hemispheres would extract distinct but overlapping features, despite identical observations.
-In addition, we expected that a bilateral architecture would be able to combine these features to improve overall accuracy.
-We chose a hierarchical image classification task where each sample belongs to both a fine and a coarse class e.g.
+Division of the brain into left and right hemispheres is a remarkably conserved feature of brains across species, suggesting the importance of the bilateral architecture for intelligence.
+The anatomy and physiology of the hemispheres have a large degree of overlap, but they are asymmetric and specialise to possess different attributes.
+A likely cause of specialisation is the hemispheric asymmetry, which includes differences in neuron density [xweems_hemispheric_2004, 10], distribution of neurotransmitters [30], firing thresholds [30] and intrahemispheric connectivity patterns [32, 26].
+One of the emergent functional differences is a specialisation for local (left hemisphere) vs global (right hemisphere) features for visual and auditory processing [13].
+In a classic experimental setup [23], a large letter is made up of smaller letters.
+Subjects are faster at detecting the small letters when presented to the left hemisphere, and the large letters when presented to the right.
+Computational models that mimic neuroanatomical and physiological asymmetries are typically subjected to standard tasks and compared to human behavioural data.
+They provide evidence that asymmetries can cause observed lateralisation of activity for specific tasks, e.g., [29, 22, 32, 37].
+The central question is usually to understand the cause of lateralisation/specialisation.
+However, none have been applied to object recognition, and none have tried to understand how differences can be exploited to improve performance on a given task.
+The aim of this project is to investigate the functional benefit of complementary hemispheres for a given task.
+Our objective is twofold.
+First, improve performance in a vision task using a bilateral architecture.
+Second, better understand the functional benefit of a bi-hemispheric architecture and understand how bilateral principles could be exploited more generally for AI/ML.
+We focused on the left hemisphere’s specialisation in local features and the right’s in global features.
+To investigate this, we chose a hierarchical image classification task where each sample belongs to both a fine and a coarse class, e.g.
 coarse: sea creature, fine: penguin, seal, shark.
+In this setup, fine labels represent local features while coarse labels represent global features.
+Our approach is to construct a bilateral network and induce specialisation in each hemisphere directly, rather than to test if it emerges as a result of architectural asymmetry.
 We modelled the hemispheres with left and right CNNs and induced specialisation using supervised training with different objectives, using either coarse or fine labels.
-The bilateral network was compared to several comparative baselines that do not have specialisation but do have the same number of learnable parameters.
+The bilateral network was compared with several baselines that do not have specialisation but do have the same number of learnable parameters.
 The differences were analysed to explore the effects of bilateral specialisation.
@@
 2 Related work
-Biologically inspired multi-network architectures
+Specialisation for local and global visual features has been attributed to differential sensitivity to spatial frequencies and explored with computational models for face perception [7, 14, 37] and a line bisection task [34, 21].
+However, other studies do not support hemispheric specialisation for particular frequency ranges, and Hsiao et al.
+[13] showed that this specialisation may emerge as a result of differences in neuron connectivity.
 Beaulieu et al.
-[20] created a dual-stream architecture called Neuromodulator Meta-learner, where one network learns to modulate the other, to enhance continual learning.
+(2020) [3] created a dual-stream architecture called Neuromodulator Meta-learner, where one network learns to modulate the other, to enhance continual learning.
 Bakhtiari et al.
-[21] created a network with parallel pathways to reproduce the functionality of dorsal (‘where’) and ventral (‘what’) pathways in an ANN trained with a single loss function.
-Li and Deza [22] discussed specialisation of a branched neural network.
+(2021) [2] created a network with parallel pathways to reproduce the functionality of dorsal (‘where’) and ventral (‘what’) pathways in an ANN trained with a single loss function.
+Li et al.
+(2021) [19] discussed specialisation of a branched neural network.
 In the most related study, Mayan et al.
-[19] tackled the same image classification task with fine and coarse (hierarchical) classes.
-The authors created a single network with two hemispheres, and trained it with supervised learning and a single loss function.
-They used hyperparameters with analogies to biological parameters in each hemisphere to encourage specialisation.
-Specialisation was achieved in each side but the bilateral network did not have an advantage over baselines.
-Ensemble models
+(2021) [20] tackled the same classification task with fine and coarse (hierarchical) labels.
+A single network with two hemispheres was trained with supervised learning and a single loss function.
+They used different hyperparameters in each hemisphere, with analogies to biological parameters, to encourage specialisation; which was achieved but the bilateral network did not have an advantage over baselines.
 Our bilateral approach can be viewed as a type of ensemble, which is a popular approach in ML [31].
@@
 Our work can be viewed as a special case of a diversified ensemble, where the ensemble architecture exploits hierarchical data labels to achieve diversification.
-Multi-domain learning
 In multi-domain learning (MDL), multiple models are trained to specialise in a particular domain [12].
@@
 In contrast, this project is focused on explicit hierarchical labels from the same distribution.
-The objective being to mimic learning different aspects of a single task with left and right hemispheres (2 separate models).
-3 Models
-We compared a bilateral model to several baselines.
+The objective being to mimic learning different aspects of a single task with left and right hemispheres.
+3 Methods
+3.1 Dataset
+We used the CIFAR-100 dataset [17], as it includes hierarchical labels that denote fine (local) and coarse (global) classes.
+There are 100 fine and 20 coarse classes.
+The dataset is split into 50,000 training and 10,000 test samples.
+Images were resized to 32x32 and we used two widely adopted data augmentation techniques, a random crop and a random horizontal flip (using inbuilt PyTorch functions) to increase diversity of the training set.
+3.2 Models and specialisation approach
+We compared a bilateral model to several baselines, shown in Figure 1 and described below.
+Details on training and evaluation in Section 3.3.
 Each model consisted of one or more feature extractor, representing hemispheres, as well as one or two classifier heads, one for each label type (fine or coarse).
@@
 We chose VGG and ResNet because they are common, well understood and amongst the best performing on many vision tasks.
+Figure 1: Bilateral architecture and baselines.
+Feature extractors are either ResNet-9 or VGG-11.
+The classifier heads are single layer fully connected networks, with output dimensions the same size as the number of classes.
+For brevity, if not specified, the network is unspecialised.
+3.2.1 Common elements – backbones and heads
 In the case of ResNet-based models, we empirically optimised the number of layers on the dataset, resulting in a 9-layer network, hereafter referred to as ResNet-9.
@@
 VGG does not have skip connections, and there are a lot more parameters due to the fully-connected layers.
-The heads consist of a single fully-connected layer, without bias or non-linearity and with output dimensions the same size as the number of classes (10 for coarse and 100 for fine).
-The models are depicted in Figure 1 and described below. Details on training and evaluation in Section 4.1.2.
-Figure 1: The range of models.
-Feature extractors are either ResNet-9 or VGG-11.
-The classifier heads are single layer fully connected networks, with output dimensions the same size as the number of classes.
-For brevity, if not specified, the network is unspecialised.
-3.1 Bilateral network - specialised
+The heads consist of a single fully-connected layer, without bias or non-linearity and with dimensions the same size as the number of classes (10 for coarse and 100 for fine).
+3.2.2 Bilateral network with specialisation
 Assembling and training the network consisted of two phases.
 In Phase 1, individual hemispheres were trained to be specialised feature extractors using supervised learning with a single classifier head, resulting in ‘Unilateral - specialised’, Figure 1a.
-The left was specialised for specific classes using fine labels, and the right for general classes with coarse labels.
+The left was specialised for local features using fine labels, and the right was specialised for global features with coarse labels.
 In Phase 2, the hemispheres (serving as feature extractors) were brought together into a broader architecture, with two initialised classifier heads, one for fine and coarse labels, resulting in ‘Bilateral - specialised’, Figure 1b, .
@@
 A single loss was calculated by adding the loss from each head.
-3.2 Baseline models
-Unilateral networks with specialisation
-We compared the bilateral network to the individual specialised hemispheres alone, on the tasks for which they were specialised, referred to as Unilateral-specialised, Figure 1a.
-Specialisation was achieved using supervised training with selective labels as described in Section 3.1.
-Unilateral network without specialisation
+3.2.3 Unilateral network with specialisation
+We compared the bilateral network with the individual specialised hemispheres alone, on the tasks for which they were specialised, referred to as Unilateral-specialised, Figure 1a.
+3.2.4 Unilateral network without specialisation
 To investigate specialisation, we compared Unilateral-specialised to a single hemisphere without specialisation, referred to as Unilateral-unspecialised, Figure 1d.
-The single hemisphere was trained on both types of labels simultaneously, by using two classifier heads and adding the loss.
+The single hemisphere was trained on both types of labels simultaneously, using two classifier heads and adding the loss.
 This baseline also helps to understand the effect of complementary loss functions; on a single hemisphere compared to the bilateral network.
-Bilateral network without specialisation
+3.2.5 Bilateral network without specialisation
 To better understand the role of specialisation in the bilateral network, we compared Bilateral-specialised to an equivalent network, with the same number of trainable parameters, but without specialisation, referred to as Bilateral-unspecialised, Figure 1c.
@@
 Like for the unilateral network and bilateral network with specialisation, the loss from each head are added together to create a single loss.
-Ensembles
+3.2.6 Ensembles
 To understand the differences between differential specialisation and conventional ensembling, we compared to a 2-model ensemble, Figure 1e.
@@
 We used two unilateral models without specialisation, trained from different seeds.
-The output of the ensemble was the average from the heads: for fine classification, the average of the fine heads, and for coarse, the average of the coarse heads.
+The output of the ensemble was the average of the heads: for fine classification, the average of the fine heads, and for coarse, the average of the coarse heads.
 We trained five models with different seeds, and then constructed five 2-model ensembles, each with a unique pair of models selected from the pool of five.
-4 Experiments
-4.1 Experimental setup
-4.1.1 Dataset
-We used the CIFAR-100 dataset [30], as it includes hierarchical labels that denote fine (specific) and coarse (general) classes.
-The dataset is split into 50,000 training and 10,000 test samples.
-Images were resized to 32x32 and we used two widely adopted data augmentation techniques, a random crop and a random horizontal flip (using the inbuilt PyTorch functions) to increase the diversity of the training set.
-4.1.2 Training and evaluation
+3.3 Training procedure
 The models were trained with supervised learning using a cross-entropy loss function.
-In the case of dual heads (one with each label type, see Section 3), the loss from each head was added to create a single loss value.
-Weights were initialised with PyTorch defaults, from a random uniform distribution $U(-\sqrt{k},\sqrt{k})$: for linear layers $k=1/in\_features$, and for the convolutional layers $k=\frac{1}{C_{in}*\Pi_{i=0}^{1}kernel\_size[i]}$.
-We used weight decay of 1.0e-5, dropout of 0.6 on the classifier heads to regularise the models, Adam [39] for optimisation, with a learning rate of 1.0e-4 and a mini-batch size of 256.
+In the case of dual heads, the loss from each head was added to create a single loss value.
+Weights were initialised with PyTorch defaults, from a random uniform distribution $U(-\sqrt{k},\sqrt{k})$: for linear layers $k=1/in\_features$, and for convolutional layers $k=\frac{1}{C_{in}*\Pi_{i=0}^{1}kernel\_size[i]}$.
+We used weight decay of 1.0e-5, dropout of 0.6 on the classifier heads to regularise the models, Adam optimiser [16], a learning rate of 1.0e-4 and mini-batch size of 256.
+These hyperparameters were empirically optimised with manual coordinate descent.
 Each network was trained for 180 epochs from 5 random seeds.
@@
 The number of epochs was sufficient to ensure that the validation accuracy had plateaued, and the final model was the one with the highest validation accuracy across epochs.
-4.1.3 Framework and computational resources
-Experiments were conducted with the PyTorch-Lightning [40] research framework, a wrapper around the PyTorch library [41].
-All models were trained and evaluated with a virtual machine with a single NVIDIA A10 GPU, 24GB RAM (through Lambda Labs https://lambdalabs.com/).
 3.4 Visualising network operation
@@
   • Scenario 4: Bilateral-specialised network is incorrect, left and right hemispheres are correct
-Gradient camera (Grad-Cam) visualisation
+3.4.1 Gradient camera (Grad-Cam) visualisation
 To understand how the specialised bilateral network exploits features from both hemispheres, we visualised gradient flow with respect to target convolutional layers [33, 5] using the Grad-Cam library [9].
@@
 In the case of the bilateral network, the gradients were averaged over both hemispheres during simultaneous prediction of both coarse and fine labels (the network had two heads).
-Feature analysis using cosine similarity
+3.4.2 Feature analysis using cosine similarity
 We investigated how left and right features are exploited by the bilateral network by analysing the relationship between representations in different parts of the network, with a focus on how the representations are transformed by the network heads.
@@
 Similarity is plotted along the following dimensions: left hemisphere, right hemisphere and a third dimension, either concatenated or bilateral.
-Additionally, univariate marginal distributions were plotted for each hemisphere.
+In addition, univariate marginal distributions were plotted for each hemisphere.
+3.5 Experimental setup
+Experiments were conducted with the PyTorch-Lightning [8] research framework, a wrapper around the PyTorch library [25].
+All models were trained and evaluated with a virtual machine with a single NVIDIA A10 GPU, 24GB RAM (through Lambda Labs https://lambdalabs.com/).
 4 Results
@@
 4.2 Visualising network operation
-The Grad-Cam visualisations are shown in Figures 4 to 7 for the scenarios described in Section 4.2.
+Grad-Cam visualisations are shown in Figures 4 to 7 for scenarios described in Section 3.4.
 In general, the features appear to be more local in the left hemisphere and more global in the right.
@@
 Furthermore, we found that it is not due to the greater number of learnable parameters, and having two hemispheres is not sufficient, but that specialisation is important, shown by the benefit of Bilateral-specialised over Bilateral.
-Surprisingly, unspecialised hemispheres (Unilateral) performed as well or better than specialised hemispheres (Unilateral-specialised) in most cases.
-Training with two objective functions (fine and coarse) led to better performance on both labels – whether that is in a bilateral or unilateral architecture.
-Ultimately, two hemispheres arranged into a conventional ensemble achieved the best results.
-A key difference with Bilateral-specialised, is that the component models were more effective to begin with.
-Another difference is the way in which the model outputs are combined – averaging of the classification vs concatenation of the feature vector.
+Surprisingly, unspecialised hemispheres (Unilateral) performed as well as or better than specialised hemispheres (Unilateral-specialised) in most cases.
+Recall that unilateral models are trained with two objective functions (fine and coarse).
+So using two training objectives led to better performance on both label types, whether that is in a bilateral or unilateral architecture.
+Using two explicit objectives could be seen as a form of specialisation, even when used simultaneously and in a single model.
+Ultimately, two hemispheres arranged in a conventional ensemble achieved the best results.
+A key difference with Bilateral-specialised is that the component models of the ensemble were more effective to begin with; benefiting from two training objectives.
+Another difference is the way that the model outputs are combined – averaging of the classification vs concatenation of the feature vector.
 Future work could explore this in more detail.
-Both the specialised bilateral network and the ensemble were able to boost performance of individual hemispheres. However, they use different mechanisms.
+Both the specialised bilateral network and the ensemble were able to boost performance of individual hemispheres.
+However, they use different mechanisms.
 The novel finding of the paper is that a bilateral architecture can exploit specialised hemispheres, and the remainder of the paper examines this principle.
-Computational cost
-Compared to a conventional single network, the bilateral architecture has approximately double the number of parameters and computational cost.
-The performance boost between unilateral and specialised bilateral models may appear modest for the additional cost.
+5.0.1 Computational cost
+The performance boost of bilateral models over a single hemisphere may appear modest relative to the additional computational cost, which could limit their practical applications.
 However, incremental gains are difficult to achieve at high levels of performance, making the  2.5-5% improvement (across fine, coarse and hemisphere architectures) consequential.
-This is the first study to demonstrate the bilateral principle, and like other new architectures, greater improvements can be expected in the future.
-How does the bilateral architecture work?
+Like other new architectures, improvements can be expected in the future.
+5.0.2 How does the bilateral architecture work?
 The Grad-Cam images reveal that the left hemisphere extracts more localised features than the right.
 Different learning objectives enable them to capture different aspects of the environment.
-Collectively the set of features is greater than one network with one objective.
-Interestingly, even though the left is explicitly trained on fine labels, the features that it extracts are helpful for coarse classes.
-The inverse is true of the right.
+Collectively, the set of features is greater than one network with one objective.
+Interestingly, even though the left is explicitly trained on fine labels, the features it extracts are helpful for coarse classes.
+The inverse is also true.
 With reference to the cosine similarity visualisation, before the network heads, the similarity of pairs of images of the same class is correlated with left and right and the similarity is no longer correlated after the network heads; which suggests that the heads implement a non-linear transformation.
@@
 In summary, specialisation creates a higher diversity of features.
-The network heads implement a type of weighted attention to left and right hemispheres selectively in a task dependent manner, improving overall class prediction.
-Relation to biology
-A limitation of our model, in terms of biological accuracy, was that inter-hemispheric connectivity was achieved at the outputs of the hemispheres.
+The network heads implement a type of weighted attention to left and right hemispheres selectively in a task-dependent manner, improving overall class prediction.
+5.0.3 Relation to biology
+A limitation in terms of biological accuracy was that interhemispheric connectivity was achieved at the outputs of the hemispheres.
 In contrast, biological hemispheres are interconnected throughout their hierarchies [4].
-Nevertheless, the heads may serve a similar purpose, albeit in a cruder way.
-Inter-hemispheric connectivity comprises a complex combination of inhibitory and excitatory projections.
-The hemisphere that is better able to represent the input is likely to have stronger activation and thus inhibit the other hemisphere [1, 2].
+However, heads may serve a similar purpose, albeit in a cruder way.
+Interhemispheric connectivity comprises a complex combination of inhibitory and excitatory projections.
+The hemisphere that is better able to represent the input is likely to have stronger activation and thus inhibit the other hemisphere [38].
 Like the heads, this too is a type of selectivity.
-Building a model with interconnected hemispheres throughout the hierarchy is a topic for future research.
-The central finding, that left and right hemispheres extracted local and non-local features, to improve performance on a given task, provides one plausible suggestion (and testable hypothesis) for the prevalence of bilateral brains in nature.
-A possible approach to investigate this idea is with transcranial magnetic stimulation (TMS) to selectively impair one hemisphere at a time on controlled tasks [46].
+Interconnected hemispheres throughout the hierarchy is a topic for future work.
+The main finding, that left and right hemispheres extracted local and non-local features to improve performance on a given task, provides one plausible reason (and testable hypothesis) for the prevalence of bilateral brains in nature.
+A possible approach to investigate this is with transcranial magnetic stimulation (TMS) to selectively impair one hemisphere at a time on controlled tasks [27].
 6 Conclusion
@@
 Simultaneously, this work shows that neuroscientific principles can provide inductive biases for novel ML architectures.
-Currently, in the field of AI/ML where scaling of existing architectures is achieving great success, it’s interesting to look at new principles on smaller architecture that could then be scaled.
-Acknowledgements
+Currently, in AI/ML where scaling of existing architectures is achieving great success, it’s interesting to look at new principles on smaller architecture that could then be scaled.
+{credits}
+6.0.1 Acknowledgements
 Thanks to Punarjay Chakravarty for helpful discussions.